How many moles of water are produced when 5.00 moles of oxygen are used?


1
Expert's answer
2021-03-11T08:46:39-0500

Solution:

The balanced chemical equation:

O2 + 2H2 → 2H2O

According to the equation above: n(O2) = n(H2O)/2

Hence,

n(H2O) = 2 × n(O2) = 2 × 5.00 mol = 10.0 mol

n(H2O) = 10.0 mol


Answer: 10.0 mol of water (H2O) are produced.